# GRACE: Generating Socially Appropriate Robot Actions Leveraging LLMs and Human Explanations

### v1

@@ Abstract
When operating in human environments, robots need to handle complex tasks while both adhering to social norms and accommodating individual preferences. For instance, based on common sense knowledge, a household robot can predict that it should avoid vacuuming during a social gathering, but it may still be uncertain whether it should vacuum before or after having guests. In such cases, integrating common-sense knowledge with human preferences, often conveyed through human explanations, is fundamental yet a challenge for existing systems. In this paper, we introduce GRACE, a novel approach addressing this while generating socially appropriate robot actions. GRACE leverages common sense knowledge from Large Language Models (LLMs), and it integrates this knowledge with human explanations through a generative network architecture. The bidirectional structure of GRACE enables robots to refine and enhance LLM predictions by utilizing human explanations and makes robots capable of generating such explanations for human-specified actions. Our experimental evaluations show that integrating human explanations boosts GRACE’s performance, where it outperforms several baselines and provides sensible explanations.
Index Terms:
Socially Appropriate Robot Actions, Large Language Models (LLMs), Explanations, Generative AI
@@ I Introduction
Robots are expected to operate in challenging real-world environments by learning complex tasks without violating social norms and human preferences. For instance, imagine a robot that is expected to perform household tasks, such as vacuum cleaning or carrying furniture. With the aid of common sense knowledge, the robot can predict that it should not perform these tasks when people are sleeping or when guests are around. However, even with such knowledge, this decision-making process gets complicated in the real world. For instance, in a home with small children, one person may emphasize safety, preferring the robot to only carry lighter items, while another might prioritize efficiency, expecting the robot to carry multiple heavy objects at once to speed up tasks. In such cases, if the robot could effectively gather insights and explanations about the factors it should base its decisions on, such as whether to prioritize efficiency, comfort, or safety, it can successfully adjust its actions according to human preferences by still adhering to social norms.
Previous research has explored socially aware robot behaviors across various domains, including household environments [1, 2, 3], navigation scenarios [4, 5, 6, 7] and service applications [8, 9, 10]. These studies have formed a critical foundation for generating robotics systems that are socially appropriate, yet in order to effectively integrate both social norms and user preferences into robotic actions, we identify the following critical open challenges: Challenge 1: How can the robot decide if is it sufficient to rely on common sense knowledge for the appropriateness of its actions and when should it seek additional insights and explanations from humans around to base its decisions on? Challenge 2: How can the robot effectively integrate these explanations into its decision-making process to ensure its actions are aligned with human expectations? Challenge 3: How can the robot generate insights or reasoning for human-specified actions, even when detailed explanations are not provided, in a way that implicitly reflects the humans’ (i.e., its user’s) underlying preferences and motivations? Our work addresses these challenges with a novel methodology that integrates robot action appropriateness with human explanations.
@@
In this paper, we present the so-called system ‘GRACE’ for its ability to generate socially appropriate robot actions. GRACE is a complete system to generate social appropriateness scores for varying robot actions (such as cleaning, serving food, vacuuming, etc) in household setups by leveraging common sense knowledge via Large Language Models (LLMs) and integrating this knowledge with human explanations through a generative deep neural network architecture. Specifically, GRACE first uses machine learning classifiers to assess the uncertainty of scenes based on human agreement, and for scenes classified as ‘certain’, it directly outputs LLM appropriateness predictions, hence addressing Challenge 1. Critically, for ‘uncertain’ cases, the conditional autoencoder structure of GRACE enhances LLM predictions by integrating them with human explanations, thus addressing Challenge 2. Additionally, the bidirectional structure of the GRACE autoencoder enables it to provide coherent explanations for human-annotated scores, thereby addressing Challenge 3. Through extensive evaluations, we demonstrate that GRACE effectively captures the relationship between human scores and their explanations, outperforms several baselines, and boosts performance by leveraging human explanations to predict social appropriateness scores for robot actions.
@@ II Related Work
While LLMs can offer a critical basis for socially aware robot behavior, integrating explanations is essential to ensure accountability [27], trust [28, 29], and the efficiency of HRI [30, 31, 32]. Despite its importance, this area of research presents significant challenges due to the black-box nature of LLMs [33], and GRACE makes a timely and novel contribution in this regard. To our knowledge, GRACE is the first full system to establish a bidirectional approach that integrates LLM-based reasoning with human explanations not only to generate socially appropriate robot actions but also to provide coherent robot explanations.
@@ III-A MannersDB and MannerDB+
To generate socially appropriate robot actions, we utilized two datasets: MANNERSDB [1] (as in [26]) and MANNERSDB+. MANNERSDB contains 750 virtual home settings with a Pepper robot and 11,050 human annotations of these scenes that scored the social appropriateness of eight robot actions (e.g., ‘vacuuming’, ‘carrying warm food’, etc.) on a scale of 1-5, along with explanations for these scores. MANNERSDB+ extends this dataset to three robots, PR2, Nao, and Pepper, across 1,000 scenes per robot. It includes social appropriateness scores for nine actions, collected from 9,238 human annotators (humans, henceforth), along with their explanations – see Figure 3 for the complete list of robot actions and example images. In both datasets, each scene includes 29 distinctive features, such as the number of people, their orientation, and whether the music is playing in the background. These scene attributes, annotated scores, and explanations make these datasets highly suitable for evaluating the objectives and capabilities of the GRACE system.
@@ III-B Categorization and Labels of Human Explanations
To integrate human explanations while generating socially appropriate robot actions, we categorized and labeled human explanations in MannersDB+ (since it has more scenes than MannersDB) by first identifying prominent explanation categories within data and then labeling each sample based on these categories using a large language model (LLM). First, a human evaluator (female, 30 years old, third author of the paper) randomly selected 100 samples from the MannersDB+ dataset and decided on the most recurring explanation categories within these random samples. The evaluator identified nine main categories (e.g., human state) and three subcategories based on two opposing (e.g., available: 1, and busy: -1) and one neutral component (neutral: 0) [^1] Identified explanation categories: human state (busy, available, neutral), human direction (facing the robot, not facing the robot, neutral), safety (safe, dangerous, neutral), robot direction (facing humans, not facing humans, neutral), working area (big, small, neutral), robot capability (capable, incapable, neutral), robot proximity (close, far, neutral), crowd (crowded, empty neutral), environmental noise (noisy, silent, neutral)..
For each category, we labeled human explanations in MannersDB+ by querying GPT4o-mini separately. We applied a two-shot prompting technique in each category by providing examples from the opposing components. To ensure the correctness of LLM labels, the same human evaluator selected another random 200 samples, manually labeled them based on pre-defined categories, and compared them with LLM-generated ones, which resulted in a high match (95.28%) and confirmed the precision of the LLM-generated labels. After obtaining the LLM labels, we had to discard human direction and environmental noise categories due to inadequate samples. Overall, for each explanation, we obtained a seven-dimensional vector, and each of these dimensions could take three different values (-1, 0, or 1), which we later normalized each dimension between 0 and 1.
@@ IV METHODOLOGY
In this section, we conceptualize the novel idea of generating socially appropriate robot actions by leveraging large language models (LLMs) and human annotator explanations. For a given scene, we first assess the uncertainty of the scene, i.e., if human annotators would agree on the scores for robot action appropriateness. To achieve this, we obtain binary labels with unsupervised clustering and train ML classifiers. This classification aims to identify when LLM predictions would be enough to predict robot action appropriateness (certain scenes, i.e., humans would agree on their scores) or when LLM predictions should be improved with explanations (uncertain scenes, i.e., humans would have disagreements). For uncertain situations, we suggest an autoencoder-based generative model that leverages human explanations to enhance the generation of socially appropriate robot actions, as well as endow the robot with the capability to generate explanations when necessary. The overall GRACE system pipeline is shown in Figure 1.
@@ IV-A Scene Clustering and Uncertainty Classification
To determine the certainty of scenes, we rely on human annotator agreement, i.e., classifying scenes based on when human annotators are most likely to give similar scores for the appropriateness of robot actions (certain scenes) or they are likely to give different ones (uncertain scenes). We formulate this task as a binary class classification problem. We first obtain the weak binary labels for each scene (showing if a scene is certain or uncertain) with unsupervised clustering and then employ ML methods to predict the scene uncertainties.
@@
After generating the weak labels for each scene $x_{i}$, these labels serve as the target variable $y_{x_{i}}\in\{0,1\}$ in a binary classification task to predict scene uncertainties. The input space of the classification tasks is obtained from scene features, i.e., for each scene, $x_{i}$, the set of scene features is $f_{x_{i}}=\{f^{1}_{x_{i}},f^{2}_{x_{i}},\dots,f^{p}_{x_{i}}\}$, where $f_{x_{i}}$ includes $p$ number of descriptive attributes such as the number of people in the scene, their relative directions, if the music is playing in the background, etc. For the prediction process, we employ varying machine learning models: logistic regression (LR), random forest (RF), support vector machine (SVM), Multi-layer perceptron (MLP), gradient boosting (GB), and k-nearest neighbors classifier (KNN) algorithms. Given a trained classifier $g$ and its model parameters $\theta$, the classifier predicts the label $\hat{y}_{x_{i}}=g(\{f^{1}_{x_{i}},f^{2}_{x_{i}},\dots,f^{p}_{x_{i}}\};\theta)$ for each scene. The training details are provided in Section V-A1.
@@ IV-C Leveraging Explanations for Action Appropriateness
$I=\begin{cases}(c_{0},S_{LLM},E_{human}),&\text{ score correction},\\ (c_{1},S_{human},E_{noise}),&\text{ explanation generation},\end{cases}$ (1)
where $c_{0}$ is the condition for score correction, and $c_{1}$ is for explanation generation. Additionally, $S_{LLM}$ shows the vector of LLM-predicted scores, whereas $S_{human}$ is the vector of human scores, which both have $n$ dimensions for different robot actions. Lastly, $E_{human}$ represents the human explanations (one-hot encoded and has $m$ dimensions denoting the different categories of human explanations, e.g., the robot’s proximity dimension is one if the human’s explanation states that the robot is far away from people, it is zero if the robot is too close to people, and 0.5 otherwise), and $E_{noise}$ is random explanations generated from noise. The target value for both conditional inputs $I_{c_{0}}$ and $I{c_{1}}$ is set as $(S_{human},E_{human})$, aiming to reconstruct human scores and explanations.
@@ V-A2 LLM Predictions
In order to obtain LLM predictions for robot actions’ social appropriateness, we experimented with Google/flan-t5-xxl, Llama3.1-8B-Instruct, Llama-3.1-70B-Instruct, GPT-4o-mini and GPT-4o language models on MannersDB+. Building on the querying method by Bowen and Harold [26], we adjusted the prompts to include the robot type and restricted the models’ responses to only output the appropriateness of the actions. For Llama 3.1 models, we further enforced this constraint by suppressing the generation of nine words, including ‘The’, ‘Answer’, and empty string.
In the experiments, the number of most probable outputs that LLM generates ($k$) was set to five to cover all the options ranging from ‘very inappropriate’ (score 1) to ‘very appropriate’ (score 5). We reported the accuracies of LLMs based on Root Mean Square Error (RMSE, lower is better), Pearson Correlation Coefficient (PCC, higher is better), and Concordance Correlation Coefficient (CCC, higher is better) to measure how closely LLM predictions matched human scores.
@@ V-B1 Uncertainty Classifications
This section presents uncertainty classification results for MannersDB and MannersDB+ by comparing ML models based on their match with the cluster-based ground truth. For MannersDB, the SVM method achieved the best performances among all models (mean values on the left side of Table I, with STD values from cross-validation ranging between 0.01 and 0.05). Oversampling improved GB performances but decreased KNN and MLP accuracies. After identifying SVM as the best model, we also applied bagging to SVM, which resulted in a slight performance increase. For MannersDB+, overall accuracies were lower, and all models performed similarly (right side of Table I, STD between 0.01 and 0.04). For models without built-in class weight balancing (GB, KNN, MLP), oversampling improved balanced accuracy and macro F1 scores but decreased the macro Precision. Bagging did not affect SVM performances. Lastly, the human baseline for both datasets (first row) showed poor performance, highlighting the difficulty of the task.
@@ V-B2 Cluster Analysis and LLM Performances
Google/flan-t5-xxl Llama 3.1 8B-Instruct Llama 3.1 70B-Instruct GPT4o-mini GPT 4o
@@
Models correcting Flan-T5 Llama 8B Llama 70B GPT4o-mini GPT 4o
@@
First, we analyzed the entire MannersDB+ and each cluster based on their aleatoric uncertainty. We compute the aleatoric uncertainty as in Jonas et al. [1] by taking the mean of log variances of human scores for each scene, i.e., $log\sigma^{2}(s_{a_{j}}^{x_{i}})$. Expectedly, the lowest value was obtained for the certain cluster (-0.59 $\pm$ 0.41), whereas it was higher for the whole data (-0.31 $\pm$ 0.56) and the highest for the uncertain cluster (0.25 $\pm$ 0.36).
Next, we show LLM prediction performances on the entire MannersDB+ and each cluster in Table II, where we compare different LLMs based on how well their predictions match with human scores. Once again, expectedly, all models showed lower RMSE values and almost always the highest correlation coefficients (PCC and CCC) for samples in the certain cluster, whereas these performances were the worst for the uncertain cluster, and the whole data was in between. Additionally, when we compare the performances of different LLM models, we see that they all reported similar RMSE values (Google/flan-t5-xxl slightly better), whereas the correlation coefficients were substantially higher for GPT 4o, meaning that GPT 4o was the best one to capture the trend in human scores.

### v2

@@ Abstract
When operating in human environments, robots need to handle complex tasks while both adhering to social norms and accommodating individual preferences. For instance, based on common sense knowledge, a household robot can predict that it should avoid vacuuming during a social gathering, but it may still be uncertain whether it should vacuum before or after having guests. In such cases, integrating common-sense knowledge with human preferences, often conveyed through human explanations, is fundamental yet a challenge for existing systems. In this paper, we introduce GRACE, a novel approach addressing this while generating socially appropriate robot actions. GRACE leverages common sense knowledge from LLMs, and it integrates this knowledge with human explanations through a generative network. The bidirectional structure of GRACE enables robots to refine and enhance LLM predictions by utilizing human explanations and makes robots capable of generating such explanations for human-specified actions. Our evaluations show that integrating human explanations boosts GRACE’s performance, where it outperforms several baselines and provides sensible explanations.
@@ I Introduction
Robots are expected to operate in challenging real-world environments by learning complex tasks without violating social norms and human preferences. For instance, imagine a robot that is expected to perform household tasks, such as vacuum cleaning or carrying furniture. With the aid of common sense knowledge, the robot can predict that it should not perform these tasks when people are sleeping or when guests are around. However, even with such knowledge, this decision-making process gets complicated in the real world. For instance, in a home with small children, one person may emphasize safety, preferring the robot to only carry lighter items, while another might prioritize efficiency, expecting the robot to carry multiple heavy objects at once to speed up tasks. In such cases, if the robot could effectively gather insights and explanations about the factors that it should base its decisions on, such as whether to prioritize efficiency, comfort, or safety, it can successfully adjust its actions to human preferences by still adhering to social norms.
Previous research has explored socially aware robot behaviours across various domains, including household environments [1, 2, 3], navigation scenarios [4, 5, 6, 7] and service applications [8, 9, 10]. These studies have formed a critical foundation for generating robotics systems that are socially appropriate, yet to effectively integrate both social norms and user preferences into robotic actions, we identify the following critical open challenges: Challenge 1: How can the robot decide if is it sufficient to rely on common sense knowledge for the appropriateness of its actions and when should it seek additional insights and explanations from humans around to base its decisions on? Challenge 2: How can the robot effectively integrate these explanations into its decision-making process to ensure its actions are aligned with human expectations? Challenge 3: How can the robot generate insights or reasoning for human-specified actions, even when detailed explanations are not provided, in a way that implicitly reflects the humans’ (i.e., its user’s) underlying preferences and motivations? Our work addresses these challenges with a novel methodology that integrates robot action appropriateness with human explanations.
@@
In this paper, we present the so-called system ‘GRACE’ for its ability to generate socially appropriate robot actions. GRACE is a complete system to generate social appropriateness scores for varying robot actions (such as cleaning, serving food, vacuuming, etc.) in household setups by leveraging common sense knowledge via Large Language Models (LLMs) and integrating this knowledge with human explanations through a generative deep neural network architecture. Specifically, GRACE first uses machine learning classifiers to assess the uncertainty of scenes based on human agreement, and for scenes classified as ‘certain’, it directly outputs LLM appropriateness predictions, hence addressing Challenge 1. Critically, for ‘uncertain’ cases, the conditional autoencoder structure of GRACE enhances LLM predictions by integrating them with human explanations (e.g., ‘there is a child next to the robot’, ‘ensure safety’), thus addressing Challenge 2. Additionally, the bidirectional structure of the GRACE autoencoder enables it to provide coherent explanations for human-annotated scores, thereby addressing Challenge 3. Through extensive evaluations, we demonstrate that GRACE effectively captures the relationship between human scores and their explanations, outperforms several baselines, and boosts performance by leveraging human explanations to predict the appropriateness of robot actions.
@@ II Related Work
While LLMs can offer a critical basis for socially aware robot behaviour, integrating explanations is essential to ensure accountability [27], trust [28, 29], and the efficiency of HRI [30, 31, 32]. Despite its importance, this area of research presents significant challenges due to the black-box nature of LLMs [33], and GRACE makes a timely and novel contribution in this regard. To our knowledge, GRACE is the first full system to establish a bidirectional approach that integrates LLM-based reasoning with human explanations not only to generate socially appropriate robot actions but also to provide coherent robot explanations.
@@ III-A MannersDB and MannerDB+
To generate socially appropriate robot actions, we utilized two datasets: MANNERSDB [1] (as in [26]) and MANNERSDB+. MANNERSDB contains 750 virtual home settings with a Pepper robot and 11,050 human annotations of these scenes that scored the social appropriateness of eight robot actions (e.g., ‘vacuuming’, ‘carrying warm food’, etc.) on a scale of 1-5, along with explanations for these scores. MANNERSDB+ extends this dataset to three robots, PR2, Nao, and Pepper, across 1,000 scenes per robot. It includes social appropriateness scores for nine actions, collected from 9,238 human annotators (humans, henceforth), along with their explanations – see Figure 3 for the complete list of robot actions and example images. In both datasets, each scene includes 29 distinctive features, such as the number of people, their orientation, and whether the music is playing in the background. These scene attributes, annotated scores, and explanations make these datasets highly suitable for evaluating the objectives and capabilities of GRACE.
@@ III-B Categorization and Labels of Human Explanations
To integrate human explanations while generating socially appropriate robot actions, we categorized and labelled human explanations in MannersDB+ (since it has more scenes than MannersDB) by first identifying prominent explanation categories within data and then labelling each sample based on these categories using an LLM. First, a human evaluator (female, 30 years old, third author of the paper) randomly selected 100 samples from the MannersDB+ dataset and decided on the most recurring explanation categories within these random samples. The evaluator identified nine main categories (e.g., human state) and three subcategories based on two opposing (e.g., available: 1, and busy: -1) and one neutral component (neutral: 0) [^1] Identified explanation categories: human state (busy, available, neutral), human direction (facing the robot, not facing the robot, neutral), safety (safe, dangerous, neutral), robot direction (facing humans, not facing humans, neutral), working area (big, small, neutral), robot capability (capable, incapable, neutral), robot proximity (close, far, neutral), crowd (crowded, empty neutral), environmental noise (noisy, silent, neutral)..
For each category, we labelled human explanations in MannersDB+ by querying GPT-4o-mini separately. We applied a two-shot prompting technique in each category by providing examples from the opposing components. To ensure the correctness of LLM labels, the same human evaluator selected another random 200 samples, manually labelled them based on pre-defined categories, and compared them with LLM-generated ones, which resulted in a high match (95.28%) and confirmed the precision of the LLM-generated labels. After obtaining the LLM labels, we had to discard human direction and environmental noise categories due to inadequate samples. Overall, for each explanation, we obtained a seven-dimensional vector, and each of these dimensions could take three different values (-1, 0, or 1), which we later normalized each dimension between 0 and 1.
@@ IV METHODOLOGY
In this section, we conceptualize the novel idea of generating socially appropriate robot actions by leveraging large language models (LLMs) and human annotator explanations. For a given scene, we first assess the uncertainty of the scene, i.e., if human annotators would agree on the scores for robot action appropriateness. To achieve this, we obtain binary labels with unsupervised clustering and train ML classifiers. This classification aims to identify when LLM predictions would be enough to predict robot action appropriateness (certain scenes, i.e., humans would agree on their scores) or when LLM predictions should be improved with explanations (uncertain scenes, i.e., humans would have disagreements). For uncertain situations, we suggest an autoencoder-based generative model that leverages human explanations to enhance the generation of socially appropriate robot actions, as well as endow the robot with the capability to generate explanations when necessary (overall pipeline in Figure 1).
@@ IV-A Scene Clustering and Uncertainty Classification
To determine the certainty of scenes, we rely on human annotator agreement, i.e., classifying scenes based on when human annotators are most likely to give similar scores for the appropriateness of robot actions (certain scenes), or they are likely to give different ones (uncertain scenes). We formulate this task as a binary class classification problem. We first obtain the weak binary labels for each scene (showing if a scene is certain or uncertain) with unsupervised clustering and then employ various classifiers to predict uncertainties.
@@
After generating the weak labels for each scene $x_{i}$, these labels serve as the target variable $y_{x_{i}}\in\{0,1\}$ in a binary classification task to predict scene uncertainties. The input space of the classification tasks is obtained from scene features, i.e., for each scene, $x_{i}$, the set of scene features is $f_{x_{i}}=\{f^{1}_{x_{i}},f^{2}_{x_{i}},\dots,f^{p}_{x_{i}}\}$, where $f_{x_{i}}$ includes $p$ number of descriptive attributes such as the number of people in the scene, their relative directions, if the music is playing in the background, etc. For the prediction process, we employ varying machine learning models: logistic regression (LR), random forest (RF), support vector machine (SVM), Multi-layer perceptron (MLP), gradient boosting (GB), and k-nearest neighbour (KNN) classifiers. Given a trained classifier $g$ and its model parameters $\theta$, the classifier predicts the label $\hat{y}_{x_{i}}=g(\{f^{1}_{x_{i}},f^{2}_{x_{i}},\dots,f^{p}_{x_{i}}\};\theta)$ for each scene. The training details are provided in Section V-A1.
@@ IV-C Leveraging Explanations for Action Appropriateness
$I=\begin{cases}(c_{0},S_{LLM},E_{human}),&\text{ score correction},\\ (c_{1},S_{human},E_{noise}),&\text{{\color[rgb]{0,0,0} expl. generation}},\end{cases}\vspace{-0.3em}$ (1)
where $c_{0}$ is the condition for score correction, and $c_{1}$ is for explanation generation. Additionally, $S_{LLM}$ shows the vector of LLM-predicted scores, whereas $S_{human}$ is the vector of human scores, which both have $n$ dimensions for different robot actions. Lastly, $E_{human}$ represents the human explanations (one-hot encoded and has $m$ dimensions denoting the different categories of human explanations, e.g., the robot’s proximity dimension is one if the human’s explanation states that the robot is far away from people, it is zero if the robot is too close to people, and 0.5 otherwise), and $E_{noise}$ is random explanations generated from noise. The target value for both conditional inputs $I_{c_{0}}$ and $I_{c_{1}}$ is set as $(S_{human},E_{human})$, aiming to reconstruct human scores and explanations.
@@ V-A2 LLM Predictions
In order to obtain LLM predictions for robot actions’ social appropriateness, we experimented with Google/flan-t5-xxl, Llama-3.1 8B-Instruct, Llama-3.1 70B-Instruct, GPT-4o-mini and GPT-4o language models on MannersDB+. Building on the querying method by Bowen and Harold [26], we adjusted the prompts to include the robot type and restricted the models’ responses to only output the appropriateness of the actions. For Llama models, we further enforced this constraint by suppressing the generation of nine words, including ‘The’, ‘Answer’, and an empty string.
In the experiments, the number of most probable outputs ($k$) was set to five to cover all the options ranging from ‘very inappropriate’ (score 1) to ‘very appropriate’ (score 5). We reported the accuracies of LLMs based on Root Mean Square Error (RMSE, lower is better), Pearson Correlation Coefficient (PCC, higher is better), and Concordance Correlation Coefficient (CCC, higher is better) to measure how closely LLM predictions matched human scores.
@@ V-B1 Uncertainty Classifications
This section presents uncertainty classification results for MannersDB and MannersDB+ by comparing ML models based on their match with the cluster-based ground truth. For MannersDB, the SVM method achieved the best performances among all models (mean values on the left side of Table I, with STD values from cross-validation ranging between 0.01 and 0.05). Oversampling improved GB performances but decreased KNN and MLP accuracies. After identifying SVM as the best model, we also applied bagging to SVM, which resulted in a slight performance increase. For MannersDB+, overall accuracies were lower, and all models performed similarly (right side of Table I, STD between 0.01 and 0.04). For models without built-in class weight balancing (GB, KNN, MLP), oversampling improved balanced accuracy and F1 scores but decreased precision. Bagging did not affect SVM performances. Lastly, the human baseline showed poor performance for both datasets, highlighting task difficulty.
@@ V-B2 Cluster Analysis and LLM Performances
Google/flan-t5-xxl Llama-3.1 8B-Instruct Llama-3.1 70B-Instruct GPT-4o-mini GPT-4o
@@
Models correcting Flan-T5 Llama 8B Llama 70B GPT-4o-mini GPT-4o
@@
First, we analyzed the entire MannersDB+ and each cluster based on their aleatoric uncertainty. To compute the uncertainty, as in Jonas et al. [1], we took the mean of log variances of human scores for each scene, i.e., $log\sigma^{2}(s_{a_{j}}^{x_{i}})$. Expectedly, the lowest value was obtained for the certain cluster (-0.59 $\pm$ 0.41), whereas it was higher for the whole data (-0.31 $\pm$ 0.56) and the highest for the uncertain cluster (0.25 $\pm$ 0.36).
Next, we show LLM prediction performances on the entire MannersDB+ and each cluster in Table II, where we compare different LLMs based on how well their predictions match with human scores. Once again, expectedly, all models showed lower RMSE values and almost always the highest correlation coefficients (PCC and CCC) for samples in the certain cluster, whereas these performances were the worst for the uncertain cluster, and the whole data was in between. Additionally, when we compare the performances of different LLM models, we see that they all reported similar RMSE values (Google/flan-t5-xxl slightly better), whereas the correlation coefficients were substantially higher for GPT-4o, making it the best at capturing the trend in human scores.